OpenStudy (anonymous):

Simplify quantity x squared plus 2 x over quantity 12 x plus 54 end quantity minus quantity 3 minus x over quantity 8 x plus 36

OpenStudy (mertsj):

\[\frac{x^2+2x}{12x+54}-\frac{3-x}{8x+36}\]

OpenStudy (mertsj):

Does that look right?

OpenStudy (mertsj):

\[\frac{x(x+2)}{6(2x+9)}-\frac{3-x}{4(2x+9)}\]

OpenStudy (mertsj):

So we see that the common denominator is 12(2x+9)

OpenStudy (anonymous):

How is it 12(2x+9) ?

OpenStudy (anonymous):

nevermind i know. lol

OpenStudy (mertsj):

\[\frac{2}{2}\times\frac{x(x+2)}{6(2x+9)}-\frac{3}{3}\times \frac{3-x}{4(2x+9)}=\]

OpenStudy (mertsj):

\[\frac{2x^2+4x}{12(2x+9)}-\frac{9-3x}{12(2x+9)}=\frac{2x^2+7x-9}{12(2x+9)}\]

OpenStudy (mertsj):

\[\frac{2x^2+7x-9}{12(2x+9)}=\frac{(2x+9)(x-1)}{12(2x+9)}=\frac{x-1}{12}\]

OpenStudy (anonymous):

Thank you so much.